# Spectral selectors and contact orderability

### v1

@@ 2.4. Contact orderability
Let us prove the statement for $O=\mathcal{L}$, the proof being similar for the other cases. Let us assume $\Lambda_{0},\Lambda_{1},\Lambda_{2}\in\mathcal{L}$ are such that $\Lambda_{0}\preceq\Lambda_{1}$ and $\Lambda_{1}\cll\Lambda_{2}$; we want to show that $\Lambda_{0}\cll\Lambda_{2}$. By smoothly concatenating isotopies, one can assume the $\Lambda_{i}$’s, $i\in\{0,1,2\}$, to be part of a non-negative isotopy $(\Lambda_{t})_{t\in[0,2]}$ that is positive for $t\in(1,2]$. Let us now adapt a construction of Fraser-Polterovich-Rosen in the proof of [35, Proposition 2.6] in order to find a positive isotopy from $\Lambda_{0}$ to $\Lambda_{2}$. Let $v:[0,2]\to\mathbb{R}$ be a smooth map such that $v(0)=v(2)=0$ with $v^{\prime}$ positive on $[0,3/2]$. For $\varepsilon>0$, the isotopy $\underline{\Lambda}^{\varepsilon}:=(\phi_{\varepsilon v(t)}^{\alpha}\Lambda_{t})$ from $\Lambda_{0}$ to $\Lambda_{2}$ is then positive for $t\in[0,3/2]$. Indeed, by Lemma 2.1, one can write $\Lambda_{t}=\psi_{t}\Lambda_{0}$ with $(\psi_{t})$ contact flow generated by a non-negative Hamiltonian map, so $(\phi^{\alpha}_{\varepsilon v(t)}\psi_{t})$ has a $\alpha$-contact Hamiltonian $H_{t}\geq\varepsilon v^{\prime}(t)$ for all $t\in[0,2]$ and so does $\underline{\Lambda}^{\varepsilon}$. Since the positivity of an isotopy is a $C^{1}$-open condition, the isotopy $\underline{\Lambda}^{\varepsilon}$ is positive on $[3/2,2]$ for $\varepsilon>0$ small enough, which brings the conclusion. The proof of the case $\Lambda_{0}\cll\Lambda_{1}$ and $\Lambda_{1}\preceq\Lambda_{2}$ is similar. ∎
@@
If $(x_{t})_{t\in[0,1]}$ is a non-negative isotopy in either $\mathcal{L}$, $-~\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, the Hamiltonian $(H_{t})$ of which is positive at some $t_{0}\in[0,1]$, then $x_{0}\cll x_{1}$.
@@
    Let $J^{1}N=T^{*}N\times\mathbb{R}$ be the $1$-jet space of some manifold $N$ endowed with the standard contact structure $\ker(\mathrm{d}z-\lambda)$, $z$ being the $\mathbb{R}$-coordinate and $\lambda$ the pull-back of the Liouville form on $T^{*}N$. This contact structure induces a contact structure on the quotient space $J^{1}N/\mathbb{Z}\partial_{z}=T^{*}N\times S^{1}$. Let $0_{N}\subset J^{1}N$ denote the zero-section and $p:J^{1}N\to J^{1}N/\mathbb{Z}\partial_{z}$ the quotient map. Then $\mathcal{L}(0_{N})$ and ~$\widetilde{\mathcal{L}}(p(0_{N}))$ are orderable when $N$ is closed (it also generalizes to compactly supported Legendrian isotopies when $N$ is open) [10, 25, 76].
@@
    Given any contact manifold, there always are many Legendrian submanifolds for which $\widetilde{\mathcal{L}}$ is unorderable. Such submanifolds can be obtain by “stabilizing” any Legendrian submanifold. For any loose Legendrian submanifold of dimension $\geq 2$, $\widetilde{\mathcal{L}}$ is unorderable [25, 50].
@@ 3.1. The Legendrian case
If $\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$) is orderable, the order spectral selectors are real-valued.
@@
The composition formula of Hamiltonian maps implies that $g_{t}\varphi_{1}$ is the time-one map of a flow generated by a positive Hamiltonian when $t$ is taken large enough. Since $\underline{\Lambda}_{t}=g_{t}\varphi_{1}\Lambda$, $\Lambda\preceq\underline{\Lambda}_{t}$ when $t$ is taken large enough. One proceeds similarly to prove that $\underline{\Lambda}_{s}\preceq\Lambda$ for $-s>0$ large enough and the proof in $\mathcal{L}$ is formally identical. ∎
@@
For every complete contact form $\alpha$ supporting $\xi$ and every $\Lambda\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$),
@@
Moreover, if $\beta=e^{f}\alpha$ for some smooth $f:M\to\mathbb{R}$, $\forall\Lambda_{0},\Lambda_{1}\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$),
@@
Let us prove the second statement regarding the comparison of ±$\ell_{\pm}^{\alpha}$ and ±$\ell_{\pm}^{\beta}$ for $\beta:=e^{f}\alpha$. The $\beta$-Hamiltonian map of $(\phi^{\alpha}_{t})$ is $\beta(R^{\alpha})=e^{f}$ while the $\beta$-Hamiltonian map of $(\phi^{\beta}_{t})$ is the constant $\equiv 1$. The comparison of Hamiltonian maps $e^{\inf f}1\leq e^{f}\leq e^{\sup f}1$ implies
@@
A contact form $\alpha$ supporting $\xi$ being fixed, if $(\Lambda_{t})_{t\in[0,1]}$ is a Legendrian isotopy of $\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$), the Hamiltonian maps of which are denoted $H_{t}:\Lambda_{t}\to\mathbb{R}$, $t\in[0,1]$, one has
@@
Given any Legendrian isotopy $(\Lambda_{t})$ joining $\Lambda_{0}$ and $\Lambda_{1}$, the Hamiltonian map of which is $(H_{t})$, the triangular identity together with Lemma 3.3 imply
@@
For every $\Lambda\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$), both ±$\ell_{\pm}(\Lambda,\underline{\Lambda})$ belong to $\mathrm{Spec}(\Lambda,\underline{\Lambda})$.
@@
Let us prove the statement in $\widetilde{\mathcal{L}}$. Let us prove that $\Pi\Lambda\cap\Pi\underline{\Lambda}_{t}=\emptyset$ and $\ell_{+}(\Lambda,\underline{\Lambda})\leq t$ implies $\ell_{+}(\Lambda,\underline{\Lambda})<t$; it would imply the result for $\ell_{+}$ and the result for $\ell_{-}$ will follow from a similar argument. Since $\Pi\Lambda\cap\Pi\underline{\Lambda}_{t}=\emptyset$, by compactness $\exists\varepsilon>0$, $\forall s\in(-\varepsilon,\varepsilon)$, $\Pi\Lambda\cap\Pi\underline{\Lambda}_{t+s}=\emptyset$. We now essentially apply the trick employed by Chernov-Nemirovski in [21, Lemma 2.2]. By compactly extending the Hamiltonian generated by the isotopy $s\mapsto\underline{\Lambda}_{t+s\varepsilon/2}$, $s\in[0,1]$, in $[0,1]\times(M\setminus\Pi\Lambda)$, one finds a compactly supported $g\in\widetilde{\mathcal{G}}$ sending $\underline{\Lambda}_{t+\varepsilon/2}$ to $\underline{\Lambda}_{t-\varepsilon/2}$ and fixing $\Lambda$ (according to the isotopy extension theorem as expressed in Lemma 2.3). By definition of $\ell_{+}$ and positivity of $\underline{\Lambda}$, $\Lambda\preceq\underline{\Lambda}_{t+\varepsilon/2}$. By invariance of $\preceq$ under the action of $\widetilde{\mathcal{G}}$, applying $g$, $\Lambda\preceq\underline{\Lambda}_{t-\varepsilon/2}$ so $\ell_{+}(\Lambda,\underline{\Lambda})\leq t-\varepsilon/2$.
@@
Given any $\Lambda\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$), if $\Lambda\succeq\underline{\Lambda}_{0}$ and $\Lambda\neq\underline{\Lambda}_{0}$, then $\ell_{+}(\Lambda,\underline{\Lambda})>0$.
@@
Let us prove the statement in $\widetilde{\mathcal{L}}$. Let $\Lambda_{0}:=\underline{\Lambda}_{0}$ and let $\alpha$ be a complete contact form supporting $\xi$. By sign invariance, it is enough to prove $\ell_{+}^{\alpha}(\Lambda,\Lambda_{0})>0$. Let us consider a non-negative path $(\Lambda_{t})$ from $\Lambda_{0}$ to $\Lambda_{1}=\Lambda$. By replacing $\Lambda$ with $\Lambda_{t}$ for some smaller $t\in(0,1]$, one can moreover assume that $\Pi\Lambda$ belongs to a Weinstein neighborhood of $\Pi\Lambda_{0}$. Therefore, one can assume that $(M,\xi)$ is an open neighborhood of the zero-section of $J^{1}\Pi\Lambda_{0}$, $\Pi\Lambda_{0}$ being identified with the zero-section and $\Pi\Lambda$ being $C^{1}$-close to $\Pi\Lambda_{0}$. Let $(H_{t})$ be a non-negative Hamiltonian, the flow $(\psi^{t})$ of which satisfies $\psi^{t}\Lambda_{0}=\Lambda_{t}$ for all $t\in[0,1]$. Since $\Pi\Lambda_{0}\neq\Pi\Lambda$, there exists $t_{0}\in[0,1)$, and a non-empty neighborhood $U\subset\Pi\Lambda_{0}$, such that $H_{t_{0}}\circ\psi^{t_{0}}(q)>0$ for $q\in U$. One can furthermore assume without loss of generality $t_{0}=0$ (by replacing $\Lambda_{0}$ with $\psi^{t_{0}}\Lambda_{0}$, $U$ with $\psi^{t_{0}}U$ etc.) so that $H_{0}(q)>0$ for all $q\in U$.
We will adapt a procedure due to Eliashberg-Polterovich (see the proof of [30, Proposition 2.1.B]). Let $(\varphi_{i})_{1\leq i\leq n}$ be a finite family of diffeomorphisms of $\Pi\Lambda_{0}$ isotopic to the identity such that $(\varphi_{i}(U))$ covers $\Pi\Lambda_{0}$ (it exists by closeness of $\Pi\Lambda_{0}$). It lifts to $J^{1}\Pi\Lambda_{0}$ as a family of contactomorphisms isotopic to identity and preserving the zero-section: associating to a diffeomorphism $\varphi$, the contactomorphism $(q,p,z)\mapsto(\varphi(q),p\circ\mathrm{d}\varphi^{-1},z)$. By cutting-off their Hamiltonian maps away from the zero section, one gets contactomorphisms $(g_{i})_{1\leq i\leq n}$ of $(M,\xi)$ fixing $\Lambda_{0}$ such that $(g_{i}(U))$ covers $\Pi\Lambda_{0}$. Let $\psi^{t}_{i}:=g_{i}\psi^{t}g_{i}^{-1}$ for $t\in[0,1]$ and $1\leq i\leq n$.
@@
Given any $\Lambda\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$), if $\ell_{-}(\Lambda,\underline{\Lambda})=\ell_{+}(\Lambda,\underline{\Lambda})=t$, then $\Lambda=\underline{\Lambda}_{t}$ (resp. for $-~\widetilde{\mathcal{L}}$ it only implies $\Pi\Lambda=\Pi\underline{\Lambda}_{t}$).
@@ 3.2. The case of contactomorphisms
Let us first deal with the case of $\mathcal{G}$. We want to apply the theorem of Tsuboi asserting that the group of $C^{1}$-contactomorphisms isotopic to identity $\mathcal{G}^{1}$ is a simple group. The group of $C^{1}$-contactomorphisms is defined by Tsuboi in [71, §2], it contains $\mathcal{G}$ and its identity component is contained in the $C^{1}$-completion of $\mathcal{G}$: $\varphi$ is a $C^{1}$-contactomorphism if it is a $C^{1}$-diffeomorphism such that $\varphi^{*}\alpha=e^{f}\alpha$ for a $C^{1}$-map $f:M\to\mathbb{R}$. As $\mathcal{G}^{1}$ is contained in the $C^{1}$-completion of $\mathcal{G}$, we endow it with the $C^{1}$-topology and the maps $c_{\pm}^{\alpha}$ naturally extend to $C^{1}$-continuous maps $\bar{c}_{\pm}^{\alpha}:\mathcal{G}^{1}\to\mathbb{R}$ (see Remark 3.11 just above). Let us define the subset $Z^{1}\subset\mathcal{G}^{1}$
@@
where $f_{g}:M\to\mathbb{R}$ is the smooth function such that $g^{*}\alpha=e^{f_{g}}\alpha$. Let us recall that $g^{*}\alpha=e^{f_{g}}\alpha$ for $f_{g}$ of class $C^{1}$ when $g\in\mathcal{G}^{1}$ so that $h\mapsto\sup|f_{h}|$ is a continuous map $\mathcal{G}\to\mathbb{R}$ that extends to $\mathcal{G}^{1}$. Therefore the identity (13) extends to those $g,\varphi$ in $\mathcal{G}^{1}$ by replacing $c^{\alpha}_{\pm}$ with its extension $\bar{c}^{\alpha}_{\pm}$. This extended identity implies that $Z^{1}$ is normal. The theorem of Tsuboi [71] then implies that $Z^{1}=\{\mathrm{id}\}$ as $\phi_{t}^{\alpha}\notin Z^{1}$ for $t\neq 0$. So $Z=\{\mathrm{id}\}$ which brings the conclusion.
In the case of $\widetilde{\mathcal{G}}$, one needs to also consider the universal cover $\widetilde{\mathcal{G}}^{1}$ of $\mathcal{G}^{1}$ (it is a genuine universal cover since $\mathcal{G}^{1}$ is locally contractible [71, §3]). We denote $\Pi^{1}$ the cover map (extending the cover map $\Pi$) and define $Z^{1}\subset\widetilde{\mathcal{G}}^{1}$ as the intersection of the zero sets of $\bar{c}_{\pm}^{\alpha}$ as above. As Before, $Z^{1}$ is a normal subgroup of $\widetilde{\mathcal{G}}^{1}$. As $\Pi^{1}$ is a surjective group morphism, $\Pi^{1}Z^{1}$ is a normal subgroup of $\mathcal{G}^{1}$ which does not contain $\phi_{t}^{\alpha}$ for $t\neq 0$ small enough. The theorem of Tsuboi then implies $\Pi^{1}Z^{1}=\{\mathrm{id}\}$, which allows us to conclude. ∎

### v2

@@ 2.4. Contact orderability
Let us prove the statement for $O=\mathcal{L}$, the proof being similar for the other cases. Let us assume $\Lambda_{0},\Lambda_{1},\Lambda_{2}\in\mathcal{L}$ are such that $\Lambda_{0}\preceq\Lambda_{1}$ and $\Lambda_{1}\cll\Lambda_{2}$; we want to show that $\Lambda_{0}\cll\Lambda_{2}$. By smoothly concatenating isotopies, one can assume the $\Lambda_{i}$’s, $i\in\{0,1,2\}$, to be part of a non-negative isotopy $(\Lambda_{t})_{t\in[0,2]}$ that is positive for $t\in(1,2]$. Let us now adapt a construction of Fraser-Polterovich-Rosen in the proof of [36, Proposition 2.6] in order to find a positive isotopy from $\Lambda_{0}$ to $\Lambda_{2}$. Let $v:[0,2]\to\mathbb{R}$ be a smooth map such that $v(0)=v(2)=0$ with $v^{\prime}$ positive on [0,3/2][0,3/2][ 0 , 3 / 2 ]. For $\varepsilon>0$, the isotopy $\underline{\Lambda}^{\varepsilon}:=(\phi_{\varepsilon v(t)}^{\alpha}\Lambda_{t})$ from $\Lambda_{0}$ to $\Lambda_{2}$ is then positive for $t\in[0,3/2]$. Indeed, by Lemma 2.3, one can write $\Lambda_{t}=\psi_{t}\Lambda_{0}$ with $(\psi_{t})$ contact flow generated by a non-negative Hamiltonian map, so $(\phi^{\alpha}_{\varepsilon v(t)}\psi_{t})$ has a $\alpha$-contact Hamiltonian $H_{t}\geq\varepsilon v^{\prime}(t)$ for all $t\in[0,2]$ and so does $\underline{\Lambda}^{\varepsilon}$. Since the positivity of an isotopy is a $C^{1}$-open condition, the isotopy $\underline{\Lambda}^{\varepsilon}$ is positive on [3/2,2][3/2,2][ 3 / 2 , 2 ] for $\varepsilon>0$ small enough, which brings the conclusion. The proof of the case $\Lambda_{0}\cll\Lambda_{1}$ and $\Lambda_{1}\preceq\Lambda_{2}$ is similar. ∎
@@
If $(x_{t})_{t\in[0,1]}$ is a non-negative isotopy in either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, the Hamiltonian $(H_{t})$ of which is positive at some $t_{0}\in[0,1]$, then $x_{0}\cll x_{1}$.
@@
    Let $J^{1}N=T^{*}N\times\mathbb{R}$ be the 111-jet space of some manifold $N$ endowed with the standard contact structure $\ker(\mathrm{d}z-\lambda)$, $z$ being the $\mathbb{R}$-coordinate and $\lambda$ the pull-back of the Liouville form on $T^{*}N$. This contact structure induces a contact structure on the quotient space $J^{1}N/\mathbb{Z}\partial_{z}=T^{*}N\times S^{1}$. Let $0_{N}\subset J^{1}N$ denote the zero-section and $p:J^{1}N\to J^{1}N/\mathbb{Z}\partial_{z}$ the quotient map. Then $\mathcal{L}(0_{N})$ and ~$\widetilde{\mathcal{L}}(p(0_{N}))$ are orderable when $N$ is closed (it also generalizes to compactly supported Legendrian isotopies when $N$ is open) [11, 26, 78].
@@
    Given any contact manifold, there always are many Legendrian submanifolds for which $\widetilde{\mathcal{L}}$ is unorderable. Such submanifolds can be obtain by “stabilizing” any Legendrian submanifold. For any loose Legendrian submanifold of dimension 2$\geq$ 2≥ 2, $\widetilde{\mathcal{L}}$ is unorderable [26, 51].
@@ 3. Order spectral selectors
In this Section we first prove Theorem 1.1 in Subsection 3.1 and then prove Theorem 1.2 in Subsection 3.2.
@@ 3.1. The Legendrian case
If $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) is orderable, the order spectral selectors are real-valued.
@@
[…]
@@
For every complete contact form $\alpha$ supporting $\xi$ and every $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$),
@@
Moreover, if $\beta=e^{f}\alpha$ for some smooth $f:M\to\mathbb{R}$, $\forall\Lambda_{0},\Lambda_{1}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$),
@@
Let us prove the second statement regarding the comparison of ±$\ell_{\pm}^{\alpha}$ and ±$\ell_{\pm}^{\beta}$ for $\beta:=e^{f}\alpha$. The $\beta$-Hamiltonian map of $(\phi^{\alpha}_{t})$ is $\beta(R^{\alpha})=e^{f}$ while the $\beta$-Hamiltonian map of $(\phi^{\beta}_{t})$ is the constant 1$\equiv$ 1≡ 1. The comparison of Hamiltonian maps $e^{\inf f}1\leq e^{f}\leq e^{\sup f}1$ implies
@@
A contact form $\alpha$ supporting $\xi$ being fixed, if $(\Lambda_{t})_{t\in[0,1]}$ is a Legendrian isotopy of $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), the Hamiltonian maps of which are denoted $H_{t}:\Lambda_{t}\to\mathbb{R}$, $t\in[0,1]$, one has
@@
Given any Legendrian isotopy $(\Lambda_{t})$ joining $\Lambda_{0}$ and $\Lambda_{1}$, the Hamiltonian map of which is $(H_{t})$, the triangle identity together with Lemma 3.3 imply
@@
For every $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), both ±$\ell_{\pm}(\Lambda,\underline{\Lambda})$ belong to $\mathrm{Spec}(\Lambda,\underline{\Lambda})$.
@@
Let us prove the statement in $\widetilde{\mathcal{L}}$. Let us prove that $\Pi\Lambda\cap\Pi\underline{\Lambda}_{t}=\emptyset$ and $\ell_{+}(\Lambda,\underline{\Lambda})\leq t$ implies $\ell_{+}(\Lambda,\underline{\Lambda})<t$; it would imply the result for $\ell_{+}$ and the result for $\ell_{-}$ will follow from a similar argument. Since $\Pi\Lambda\cap\Pi\underline{\Lambda}_{t}=\emptyset$, by compactness $\exists\varepsilon>0$, $\forall s\in(-\varepsilon,\varepsilon)$, $\Pi\Lambda\cap\Pi\underline{\Lambda}_{t+s}=\emptyset$. We now essentially apply the trick employed by Chernov-Nemirovski in [22, Lemma 2.2]. By compactly extending the Hamiltonian generated by the isotopy $s\mapsto\underline{\Lambda}_{(t+\varepsilon/2)-s\varepsilon}$, $s\in[0,1]$, in $[0,1]\times(M\setminus\Pi\Lambda)$, one finds a compactly supported $g\in\widetilde{\mathcal{G}}$ sending $\underline{\Lambda}_{t+\varepsilon/2}$ to $\underline{\Lambda}_{t-\varepsilon/2}$ and fixing $\Lambda$ (according to the isotopy extension theorem as expressed in Lemma 2.5). By definition of $\ell_{+}$ and positivity of $\underline{\Lambda}$, $\Lambda\preceq\underline{\Lambda}_{t+\varepsilon/2}$. By invariance of $\preceq$ under the action of $\widetilde{\mathcal{G}}$ (see Proposition 2.9), applying $g$, $\Lambda\preceq\underline{\Lambda}_{t-\varepsilon/2}$ so $\ell_{+}(\Lambda,\underline{\Lambda})\leq t-\varepsilon/2$.
@@
Given any $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), if $\Lambda\succeq\underline{\Lambda}_{0}$ and $\Lambda\neq\underline{\Lambda}_{0}$, then $\ell_{+}(\Lambda,\underline{\Lambda})>0$.
@@
Let us prove the statement in $\widetilde{\mathcal{L}}$. Let $\Lambda_{0}:=\underline{\Lambda}_{0}$ and let $\alpha$ be a complete contact form supporting $\xi$. By sign invariance, it is enough to prove $\ell_{+}^{\alpha}(\Lambda,\Lambda_{0})>0$. Let us consider a non-negative path $(\Lambda_{t})$ from $\Lambda_{0}$ to $\Lambda_{1}=\Lambda$. Since $(\Lambda_{t})$ is non-negative $\ell_{+}(\Lambda,\Lambda_{0})\geq\ell_{+}(\Lambda_{t},\Lambda_{0})$ for all $t\in[0,1]$. Thus replacing $\Lambda$ by $\Lambda_{t}$ and taking $t\in(0,1]$ to be small enough one can moreover assume that $\Pi\Lambda$ belongs to a Weinstein neighborhood of $\Pi\Lambda_{0}$. Therefore, one can assume that $(M,\xi)$ is an open neighborhood of the zero-section of $J^{1}\Pi\Lambda_{0}$, $\Pi\Lambda_{0}$ being identified with the zero-section and $\Pi\Lambda$ being $C^{1}$-close to $\Pi\Lambda_{0}$. Let $(H_{t})$ be a non-negative Hamiltonian, the flow $(\psi^{t})$ of which satisfies $\psi^{t}\Lambda_{0}=\Lambda_{t}$ for all $t\in[0,1]$. Since $\Pi\Lambda_{0}\neq\Pi\Lambda$, there exists $t_{0}\in[0,1)$, and a non-empty neighborhood $U\subset\Pi\Lambda_{0}$, such that $H_{t_{0}}\circ\psi^{t_{0}}(q)>0$ for $q\in U$. One can furthermore assume without loss of generality $t_{0}=0$ (by replacing $\Lambda_{0}$ with $\psi^{t_{0}}\Lambda_{0}$, $U$ with $\psi^{t_{0}}U$ etc.) so that $H_{0}(q)>0$ for all $q\in U$.
We will adapt a procedure due to Eliashberg-Polterovich (see the proof of [31, Proposition 2.1.B]). Let $(\varphi_{i})_{1\leq i\leq n}$ be a finite family of diffeomorphisms of $\Pi\Lambda_{0}$ isotopic to the identity such that $(\varphi_{i}(U))$ covers $\Pi\Lambda_{0}$ (it exists by closeness of $\Pi\Lambda_{0}$). It lifts to $J^{1}\Pi\Lambda_{0}$ as a family of contactomorphisms isotopic to identity and preserving the zero-section: associating to a diffeomorphism $\varphi$, the contactomorphism $(q,p,z)\mapsto(\varphi(q),p\circ\mathrm{d}\varphi^{-1},z)$. By cutting-off their Hamiltonian maps away from the zero section, one gets contactomorphisms $(g_{i})_{1\leq i\leq n}$ of $(M,\xi)$ fixing $\Lambda_{0}$ such that $(g_{i}(U))$ covers $\Pi\Lambda_{0}$. Let $\psi^{t}_{i}:=g_{i}\psi^{t}g_{i}^{-1}$ for $t\in[0,1]$ and $1\leq i\leq n$.
@@
Given any $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), if $\ell_{-}(\Lambda,\underline{\Lambda})=\ell_{+}(\Lambda,\underline{\Lambda})=t$, then $\Lambda=\underline{\Lambda}_{t}$ (resp. for $\widetilde{\mathcal{L}}$ it only implies $\Pi\Lambda=\Pi\underline{\Lambda}_{t}$).
@@ 3.2. The case of contactomorphisms
Let us first deal with the case of $\mathcal{G}$. We want to apply the theorem of Tsuboi asserting that the group of $C^{1}$-contactomorphisms isotopic to identity $\mathcal{G}^{1}$ is a simple group. The group of $C^{1}$-contactomorphisms is defined by Tsuboi in [73, §2], it contains $\mathcal{G}$ and its identity component is contained in the $C^{1}$-completion of $\mathcal{G}$: $\varphi$ is a $C^{1}$-contactomorphism if it is a $C^{1}$-diffeomorphism such that $\varphi^{*}\alpha=e^{f}\alpha$ for a $C^{1}$-map $f:M\to\mathbb{R}$. As $\mathcal{G}^{1}$ is contained in the $C^{1}$-completion of $\mathcal{G}$, we endow it with the $C^{1}$-topology and the maps $c_{\pm}^{\alpha}$ naturally extend to $C^{1}$-continuous maps $\bar{c}_{\pm}^{\alpha}:\mathcal{G}^{1}\to\mathbb{R}$ (see Remark 3.11 just above). Let us define the subset $Z^{1}\subset\mathcal{G}^{1}$
@@
where $f_{g}:M\to\mathbb{R}$ is the smooth function such that $g^{*}\alpha=e^{f_{g}}\alpha$. Let us recall that $g^{*}\alpha=e^{f_{g}}\alpha$ for $f_{g}$ of class $C^{1}$ when $g\in\mathcal{G}^{1}$ so that $h\mapsto\sup|f_{h}|$ is a continuous map $\mathcal{G}\to\mathbb{R}$ that extends to $\mathcal{G}^{1}$. Therefore the identity (14) extends to those $g,\varphi$ in $\mathcal{G}^{1}$ by replacing $c^{\alpha}_{\pm}$ with its extension $\bar{c}^{\alpha}_{\pm}$. This extended identity implies that $Z^{1}$ is normal. The theorem of Tsuboi [73] then implies that $Z^{1}=\{\mathrm{id}\}$ as $\phi_{t}^{\alpha}\notin Z^{1}$ for $t\neq 0$ by the normalization property of Theorem 1.2. So $Z=\{\mathrm{id}\}$ which brings the conclusion.
In the case of $\widetilde{\mathcal{G}}$, one needs to also consider the universal cover $\widetilde{\mathcal{G}}^{1}$ of $\mathcal{G}^{1}$ (it is a genuine universal cover since $\mathcal{G}^{1}$ is locally contractible [73, §3]). We denote $\Pi^{1}$ the cover map (extending the cover map $\Pi$) and define $Z^{1}\subset\widetilde{\mathcal{G}}^{1}$ as the intersection of the zero sets of $\bar{c}_{\pm}^{\alpha}$ as above. As Before, $Z^{1}$ is a normal subgroup of $\widetilde{\mathcal{G}}^{1}$. As $\Pi^{1}$ is a surjective group morphism, $\Pi^{1}Z^{1}$ is a normal subgroup of $\mathcal{G}^{1}$ which does not contain $\phi_{t}^{\alpha}$ for $t\neq 0$ small enough. The theorem of Tsuboi then implies $\Pi^{1}Z^{1}=\{\mathrm{id}\}$, which allows us to conclude. ∎